# Global weak solutions for the inverse mean curvature flow in the Heisenberg group

### v2

@@ 3. Construction of barriers
Let $1<p<Q$ and $\bar{r}>0$ be fixed and for any $\varepsilon>0$ and $\bar{g}\in\mathbb{H}^{n}$ let $v=v^{\varepsilon}_{p}\in C^{1}(B^{\varepsilon}_{9\bar{r}}(\bar{g}))$ be a nonnegative $p$-harmonic function w.r. to the metric $|\cdot|^{2}_{\varepsilon}$. There exists a constant $C_{H}\geq 1$ depending only on $p,Q$ and $\bar{r}$ such that
(4.16) $\underset{B^{\varepsilon}_{r}(\bar{g})}{\rm{ess}\sup}\,\,v\leq C_{H}\,\underset{B^{\varepsilon}_{r}(\bar{g})}{\rm{ess}\inf}\,\,v\qquad\hbox{for any }0<r\leq\frac{1}{2}\bar{r}\,.$
@@
We first notice that $v>0$ in $B^{\varepsilon}_{9\bar{r}}(\bar{g})$ by the strong maximum principle, unless $v\equiv 0$ and (4.16) holds trivially. Now we derive a Caccioppoli-type inequality for arbitrary powers of $v$. Thus, for fixed $\varepsilon>0$, $\beta\in\mathbb{R}\setminus\{-1/p\}$ and for nonnegative $d_{\varepsilon}$-Lipschitz function $\zeta$ compactly supported in $B^{\varepsilon}_{9\bar{r}}(\bar{g})$ to be specified later, we test the equation for $v$ with $\zeta^{p}v^{1+\beta p}$, hence integration by parts, Cauchy-Schwarz and Young inequalities yield
$\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\zeta^{p}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}v^{\beta p}=-\frac{1}{1+\beta p}\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}|\nabla_{\varepsilon}v|_{\varepsilon}^{p-2}v^{1+\beta p}\langle\nabla_{\varepsilon}v,\nabla_{\varepsilon}\zeta^{p}\rangle_{\varepsilon}$
$\leq\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\left(\zeta^{p-1}|\nabla_{\varepsilon}v|_{\varepsilon}^{p-1}v^{\beta(p-1)}\right)\left(\frac{p}{|1+\beta p|}v^{1+\beta}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)$
$\leq\left(1-\frac{1}{p}\right)\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\zeta^{p}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}v^{\beta p}+\frac{1}{p}\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\left(\frac{p}{|1+\beta p|}\right)^{p}v^{(1+\beta)p}|\nabla_{\varepsilon}\zeta|^{p}_{\varepsilon}\,.$
Thus, for $\beta=-1$ the previous inequality in lerms of $\log v$ becomes
(4.17) $\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\zeta^{p}|\nabla_{\varepsilon}\log v|_{\varepsilon}^{p}\leq\left(\frac{p}{p-1}\right)^{p}\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}|\nabla_{\varepsilon}\zeta|^{p}_{\varepsilon}\,,$
@@
$\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\zeta^{p}|\nabla_{\varepsilon}v^{1+\beta}|_{\varepsilon}^{p}\leq\left(\frac{|1+\beta|p}{|1+\beta p|}\right)^{p}\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\left(v^{(1+\beta)}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)^{p}\,,$
which combined with Leibniz rule and a convexity inequality gives
(4.18) $\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}|\nabla_{\varepsilon}\left(\zeta v^{1+\beta}\right)|_{\varepsilon}^{p}\leq 2^{p-1}\left[\left(\frac{|1+\beta|p}{|1+\beta p|}\right)^{p}+1\right]\int_{B^{\varepsilon}_{9\bar{r}}(\bar{g})}\left(v^{(1+\beta)}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)^{p}\,.$
Next, we specialize inequalities (4.17) and (4.18) by chosing test functions adapted to the metric $d_{\varepsilon}$. Instead of resorting to the general theory of spaces of homogenous type as in, e.g., [9, Section 7], here we argue as in Lemma 2.2. Notice that $|\nabla_{\varepsilon}d_{\varepsilon}(\cdot,\bar{g})|_{\varepsilon}\leq 1$ on $\mathbb{H}^{n}$ for any $\bar{g}\in\mathbb{H}^{n}$ and for each $\varepsilon>0$ because such function is $1$-Lipschitz w.r.to $d_{\varepsilon}$ just by triangle inequality. Thus, if $0<\rho_{1}<\rho_{2}\leq 2\rho_{1}$ and $h:[0,\infty)\to[0,1]$ is a Lipschitz and piecewhise linear function, with $\{h\equiv 1\}=[0,\rho_{1}]$, $\{h\equiv 0\}=[\frac{\rho_{1}+\rho_{2}}{2},\infty)$, and $h$ is linear otherwise, then for any $\bar{g}\in\mathbb{H}^{n}$ the function $\zeta(g)=h(d_{\varepsilon}(g,\bar{g}))$ is Lipschitz on $\mathbb{H}^{n}$, with $\zeta\equiv 1$ on $B^{\varepsilon}_{\rho_{1}}(\bar{g})$, $spt\,\zeta\subset B^{\varepsilon}_{\rho_{2}}(\bar{g})$ and $|\nabla_{\varepsilon}\zeta|_{\varepsilon}\leq\frac{2}{\rho_{2}-\rho_{1}}\,$.
Thus, for $\hat{g}\in B^{\varepsilon}_{\bar{r}}(\bar{g})$ and $0<\rho\leq 2\bar{r}$ we have $B^{\varepsilon}_{3\rho}(\hat{g})\subset B^{\varepsilon}_{7\bar{r}}(\bar{g})$ and $B^{\varepsilon}_{4\rho}(\hat{g})\subset B^{\varepsilon}_{9\bar{r}}(\bar{g})$, hence for $(\rho_{1},\rho_{2})=(3\rho,4\rho)$ in view of the properties of $\zeta$ inequality (4.17) together with the doubling property (2.2) and the uniform Poincaré inequality (2.3) yield for $w=\log v$
@@
where $C>0$ depends only on $p$ and $Q$. Then, setting $\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})=B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{\bar{r}}(\bar{g})$ and applying inequality (2.8) we easily infer from (4.19) and the triangle inequality that for $\hat{g}\in B^{\varepsilon}_{\bar{r}}(\bar{g})$ and $0<\rho\leq 2\bar{r}$ we have
(4.20) $\frac{1}{|\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})|}\int_{\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})}|w-w_{\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})}|$ $\leq C\frac{1}{|B^{\varepsilon}_{\rho}(\hat{g})|^{2}}\int_{B^{\varepsilon}_{\rho}(\hat{g})}\int_{B^{\varepsilon}_{\rho}(\hat{g})}|w(g)-w(g^{\prime})|$
@@
where $C>0$ depends only on $p$ and $Q$. In particular, the last estimate shows that $w\in BMO(B^{\varepsilon}_{\bar{r}}(\bar{g}),d_{\varepsilon},|\cdot|)$ as the ball $\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})\subset B^{\varepsilon}_{\bar{r}}(\bar{g})$ is an arbitrary ball in the metric space $(B^{\varepsilon}_{\bar{r}}(\bar{g}),d_{\varepsilon})$.
Note that for all $\varepsilon>0$ the spaces $(B^{\varepsilon}_{\bar{r}}(\bar{g}),d_{\varepsilon},|\cdot|)$ are of homogeneous type, the doubling property for them holds uniformly w.r.to $\varepsilon>0$ due to (2.2) together with (2.8). Then, it follows from the bound of the BMO norms (4.20) and the abstract John-Nirenberg inequality (see, e.g, [5, Theorem 1]) that there exists $\hat{p}>0$ and for each $\bar{p}\in(0,\hat{p})$ there exists a corresponding $C=C(\bar{p})>0$ both dependent on $p$ and $Q$, possibly depending on $\bar{r}$ but independent of $\varepsilon$, such that $\int_{B^{\varepsilon}_{r}(\bar{g})}exp\left(\bar{p}|w-w_{B^{\varepsilon}_{r}(\bar{g})}|\right)\leq C|B^{\varepsilon}_{r}(\bar{g})|$ for any $0<r\leq\bar{r}$. Then, the identity $w=\log v$ and $e^{\pm\bar{p}\log v}=v^{\pm\bar{p}}$ together with exponential integrabiliy readily yield for any $0<r\leq\bar{r}$ the estimate
(4.21) $\left(\frac{1}{|B^{\varepsilon}_{r}(\bar{g})|}\int_{B^{\varepsilon}_{r}(\bar{g})}v^{\bar{p}}\right)^{1/\bar{p}}\leq C\left(\frac{1}{|B^{\varepsilon}_{r}(\bar{g})|}\int_{B^{\varepsilon}_{r}(\bar{g})}v^{-\bar{p}}\right)^{-1/\bar{p}}\,,$
where $C>0$ depends only on $\bar{p}$, $p$, $Q$ and $\bar{r}$.
@@
In order to proceed in the proof we go back to (4.18) and use the localized Sobolev inequality for fixed $0<r\leq\bar{r}$. Taking the cut-off function $\zeta$ as above, since $spt\,\zeta\subset B_{\rho_{2}}^{\varepsilon}(\bar{g})\subseteq B^{\varepsilon}_{r}(\bar{g})$, from inequality (2.5) and (2.2) we infer for any $\beta\in\mathbb{R}\setminus\{-1,-1/p\}$ and $r/4\leq\rho_{1}<\rho_{2}\leq r$ that
(4.22) $\left(\frac{1}{|B^{\varepsilon}_{\rho_{1}}(\bar{g})|}\int_{B^{\varepsilon}_{\rho_{1}}(\bar{g})}\left(v^{(1+\beta)}\right)^{\frac{pQ}{(Q-p)}}\right)^{\frac{Q-p}{pQ|1+\beta|}}$
$\leq\left(\frac{C\rho_{2}}{\rho_{2}-\rho_{1}}\right)^{\frac{1}{|1+\beta|}}\left[\left(\frac{|1+\beta|p}{|1+\beta p|}\right)^{p}+1\right]^{\frac{1}{p|\beta+1|}}\left(\frac{1}{|B^{\varepsilon}_{\rho_{2}(\bar{g})}|}\int_{B^{\varepsilon}_{\rho_{2}}(\bar{g})}\left(v^{(1+\beta)}\right)^{p}\right)^{\frac{1}{p|\beta+1|}},$
where the constant $C>0$ depends only on $p$ and $Q$.
Next we perform Moser iteration with exponents $\{\beta_{i}\}_{i\geq 0}$ where $\beta_{0}\in(-1,-1/p)\cup(-1/p,\infty)$ is to be chosen later and $\beta_{i+1}=p/(Q-p)+\beta_{i}Q/(Q-p)$, so that $\{\beta_{i}\}\subset(-1,\infty)$ and it is increasing. Note that if $s=p(\beta+1)$ then $\beta>-1$ iff $s>0$ and $\beta\neq-1/p$ iff $s\neq p-1$. Hence, setting for brevity $s_{i}=p(\beta_{i}+1)$, we have that $s_{i+1}=s_{i}Q/(Q-p)=s_{0}\frac{Q^{i+1}}{(Q-p)^{i+1}}\to+\infty$ as $i\to\infty$, so that the sequence $\{\beta_{i}\}$ is eventually nonnegative and increasing to +$\infty$+ ∞ as $\beta_{i+1}\geq C\frac{Q^{i}}{(Q-1)^{i}}$ for any $i\geq 0$ large enough. Moreover, for such positive $\beta=\beta_{i}$ one has $\left(\frac{|1+\beta|p}{|1+\beta p|}\right)^{p}+1\leq 2p^{p}$ Thus, we can select $s_{0}=\bar{p}\in(0,\hat{p})$ such that $s_{i}\neq p-1$ for all $i\geq 0$ (i.e., $\beta_{i}\in(-1,-1/p)\cup(-1/p,\infty)$ for all $i\geq 0$).
@@
Thus, for fixed $0<r\leq\bar{r}$ we repeatedly use (4.22) with the choice $(\rho_{1},\rho_{2},\beta)=(\tilde{\rho}_{i+1},\tilde{\rho}_{i},\beta_{i})$, where $\tilde{\rho}_{i}=r(\frac{1}{2}+2^{-i-1})$ for $i\geq 0$, so that inequality(4.8) rewrites for all $i\geq 0$ as
(4.23) $\left(\frac{1}{|B^{\varepsilon}_{\tilde{\rho}_{i+1}}(\bar{g})|}\int_{B^{\varepsilon}_{\tilde{\rho}_{i+1}}(\bar{g})}v^{s_{i+1}}\right)^{\frac{1}{|s_{i+1}|}}\leq C^{\frac{i+1}{|s_{i}|}}\left(\frac{1}{|B^{\varepsilon}_{\tilde{\rho}_{i}}(\bar{g})|}\int_{B_{\tilde{\rho}_{i}}^{\varepsilon}(\bar{g})}v^{s_{i}}\right)^{\frac{1}{|s_{i}|}}\,\,,$
for a constant $C>0$ depending only on $p$ and $Q$.
@@
(4.24) $\|v\|_{L^{\infty}(B^{\varepsilon}_{r/2}(\bar{g}))}\leq C^{\prime}\left(\frac{1}{|B_{r}^{\varepsilon}(\bar{g})|}\int_{B_{r}^{\varepsilon}(\bar{g})}v^{\bar{p}}\right)^{\frac{1}{\bar{p}}}\,,$
where $C^{\prime}=\Pi_{i=0}^{\infty}C^{\frac{i+1}{|s_{i}|}}<\infty$ depends only on $p$ and $Q$ (notice that $s_{i}=s_{0}\frac{Q^{i}}{(Q-p)^{i}}$ for any $i\geq 0$, whence $\sum_{i=0}^{\infty}\frac{i+1}{s_{i}}<\infty$ and in turn the convergence of the infinite product follows).
On the other hand, for fixed $0<r\leq\bar{r}$ we perform again Moser iteration to show that
(4.25) $\left(\underset{B^{\varepsilon}_{r/2}(\bar{g})}{\rm{ess}\inf}\,\,v\right)^{-1}=\left\|v^{-1}\right\|_{L^{\infty}(B^{\varepsilon}_{r/2})}\leq\widetilde{C}\left(\frac{1}{|B^{\varepsilon}_{r}(\bar{g})|}\int_{B^{\varepsilon}_{r}(\bar{g})}v^{-\bar{p}}\right)^{1/\bar{p}}\,,$
where $\widetilde{C}>0$ depends only on $\bar{p}$, $Q$ and $\bar{r}$. Indeed, note that for $\{\beta_{i}\}_{i\geq 0}$ and $\{s_{i}\}_{i\geq 0}$ as above and $\beta_{0}<-1$ one has $\beta_{i}<-1$ and in turn $\beta_{i+1}<\beta_{i}$ for all $i\geq 0$. On the other hand for $s_{0}<0$ one has both $s_{i}=s_{0}\left(\frac{Q}{Q-p}\right)^{i}\to-\infty$ and in turn $\beta_{i}\to-\infty$ as $i\to\infty$. Then, we can chose $\beta_{0}<-1$ such that $s_{0}=p(1+\beta_{0})=-\bar{p}$ and as above we can repeatedly use (4.22) with the same choice $(\rho_{1},\rho_{2},\beta)=(\tilde{\rho}_{i+1},\tilde{\rho}_{i},\beta_{i})$ and where still $\tilde{\rho}_{i}=r(\frac{1}{2}+2^{-i-1})$ for $i\geq 0$ to get once more (4.23) for all $i\geq 0$. As in the proof of (4.24), iterating (4.23) for these new set of exponents inequality (4.25) follows easily as $i\to\infty$
Finally, combining inequalities (4.24), (4.21), and (4.25) the desired inequality (4.16) follows with $C_{H}>0$ depending only on $p$, $Q$ and $\bar{r}$. ∎

### v3

@@ 3. Construction of barriers
Let $1<p<Q$, $\varepsilon>0$ and $\bar{g}\in\mathbb{H}^{n}$. For any $\bar{r}\geq 2\varepsilon$ let $v=v^{\varepsilon}_{p}\in C^{1}(B^{\varepsilon}_{9\bar{r}}(\bar{g}))$ be a nonnegative $p$-harmonic function w.r. to the metric $|\cdot|^{2}_{\varepsilon}$. There exist a constant $C>1$ and $C^{\prime}>0$ depending only on $Q$ such that
(4.16) $\underset{B^{\varepsilon}_{r}(\bar{g})}{\rm{ess}\sup}\,\,v\leq\left(Cp^{*}\right)^{\frac{C^{\prime}}{p-1}}\,\underset{B^{\varepsilon}_{r}(\bar{g})}{\rm{ess}\inf}\,\,v\qquad\hbox{for any }0<r\leq\frac{1}{2}\bar{r}\,.$
@@
It is clearly enough to prove the claim assuming $r=\frac{1}{2}\bar{r}$ to get immediately (4.16) at any smaller scale with the same constant. Moreover, by left invariance of the orthonormal frame $\{X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n},T_{\varepsilon}\}$ and the related equivariance under left translation of the corresponding $p$-Laplacian, it follows that inequality (4.16) is translation invariant and it is enough to prove it for $\bar{g}=0$, denoting in the sequel $B^{\varepsilon}_{r}=B^{\varepsilon}_{r}(0)$ for brevity. Finally, once $\bar{g}=0$ we recall that for $\widetilde{\varepsilon}:=\frac{2\varepsilon}{\bar{r}}\in(0,1]$ by group dilation we have $\delta_{\bar{r}/2}(B_{18}^{\widetilde{\varepsilon}})=B^{\varepsilon}_{9\bar{r}}$ therefore $\tilde{v}=v\circ\delta_{\bar{r}/2}$ is $p$-harmonic w.r. to $|\cdot|_{\widetilde{\varepsilon}}$ in $B_{18}^{\widetilde{\varepsilon}}$. As a consequence in what follows without loss of generality we will assume throughout the proof that $\varepsilon\in(0,1]$ and $\bar{r}=2=2r$ so that in proving (4.16) in this (rescaled) ball we will obtain constant independent of $\bar{r}$.
@@
We first notice that $v>0$ in $B^{\varepsilon}_{9\bar{r}}$ by the strong maximum principle, unless $v\equiv 0$ and (4.16) holds trivially. Now we derive a Caccioppoli-type inequality for arbitrary powers of $v$. Thus, for fixed $\varepsilon>0$, $\beta\in\mathbb{R}\setminus\{-1/p\}$ and for nonnegative $d_{\varepsilon}$-Lipschitz function $\zeta$ compactly supported in $B^{\varepsilon}_{9\bar{r}}$ to be specified later, we test the equation for $v$ with $\zeta^{p}v^{1+\beta p}$, hence integration by parts, Cauchy-Schwarz and Young inequalities yield
$\int_{B^{\varepsilon}_{9\bar{r}}}\zeta^{p}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}v^{\beta p}=-\frac{1}{1+\beta p}\int_{B^{\varepsilon}_{9\bar{r}}}|\nabla_{\varepsilon}v|_{\varepsilon}^{p-2}v^{1+\beta p}\langle\nabla_{\varepsilon}v,\nabla_{\varepsilon}\zeta^{p}\rangle_{\varepsilon}$
$\leq\int_{B^{\varepsilon}_{9\bar{r}}}\left(\zeta^{p-1}|\nabla_{\varepsilon}v|_{\varepsilon}^{p-1}v^{\beta(p-1)}\right)\left(\frac{p}{|1+\beta p|}v^{1+\beta}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)$
$\leq\left(1-\frac{1}{p}\right)\int_{B^{\varepsilon}_{9\bar{r}}}\zeta^{p}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}v^{\beta p}+\frac{1}{p}\int_{B^{\varepsilon}_{9\bar{r}}}\left(\frac{p}{|1+\beta p|}\right)^{p}v^{(1+\beta)p}|\nabla_{\varepsilon}\zeta|^{p}_{\varepsilon}\,.$
Thus, for $\beta=-1$ the previous inequality in terms of $w=(1-p)\log v$ becomes
(4.17) $\int_{B^{\varepsilon}_{9\bar{r}}}\zeta^{p}|\nabla_{\varepsilon}w|_{\varepsilon}^{p}\leq Q^{p}\int_{B^{\varepsilon}_{9\bar{r}}}|\nabla_{\varepsilon}\zeta|^{p}_{\varepsilon}\,$,
@@
(4.18) $\int_{B^{\varepsilon}_{9\bar{r}}}\zeta^{p}\left|\nabla_{\varepsilon}v^{1+\beta}\right|_{\varepsilon}^{p}\leq\left(\frac{|1+\beta|p}{|1+\beta p|}\right)^{p}\int_{B^{\varepsilon}_{9\bar{r}}}\left(v^{(1+\beta)}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)^{p}\,$,
so that for $\beta=-1+\kappa(1-1/p)$ and $\kappa\not\in\{0,1\}$ we obtain
$\int_{B^{\varepsilon}_{9\bar{r}}}\zeta^{p}\left|\nabla_{\varepsilon}v^{\kappa(1-1/p)}\right|_{\varepsilon}^{p}\leq\left(\frac{|\kappa|}{|\kappa-1|}\right)^{p}\int_{B^{\varepsilon}_{9\bar{r}}}\left(v^{\kappa(1-1/p))}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)^{p}\,$,
which combined with Leibniz rule and triangle inequality, together with convexity of $s\to s^{p}$ and subadditivity of $s\to s^{1/p}$ for $s\geq 0$, gives
(4.19) $\left(\int_{B^{\varepsilon}_{9\bar{r}}}\left|\nabla_{\varepsilon}\left(\zeta v^{\kappa(1-1/p)}\right)\right|_{\varepsilon}^{p}\right)^{1/p}\leq 4\left[\frac{1}{|\kappa-1|}+1\right]\left(\int_{B^{\varepsilon}_{9\bar{r}}}\left(v^{\kappa(1-1/p)}|\nabla_{\varepsilon}\zeta|_{\varepsilon}\right)^{p}\right)^{1/p}\,.$
[…]
Thus, for $\hat{g}\in B^{\varepsilon}_{\bar{r}}$ and $0<\rho\leq 2\bar{r}$ we have $B^{\varepsilon}_{3\rho}(\hat{g})\subset B^{\varepsilon}_{7\bar{r}}$ and $B^{\varepsilon}_{4\rho}(\hat{g})\subset B^{\varepsilon}_{9\bar{r}}$, hence for $(\rho_{1},\rho_{2})=(3\rho,4\rho)$ in view of the properties of $\zeta$ inequality (4.17) together with the doubling property (2.4), the uniform Poincaré inequality (2.5) and Holder inequality yield for $w=(1-p)\log v$
@@
where $C>0$ depends only on $Q$. Then, setting $\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})=B^{\varepsilon}_{\rho}(\hat{g})\cap B^{\varepsilon}_{\bar{r}}$ and applying inequality (2.10) we easily infer from (4.20) and the triangle inequality that for $\hat{g}\in B^{\varepsilon}_{\bar{r}}$ and $0<\rho\leq 2\bar{r}$ we have
[…]
@@
where $C>0$ depends only on $Q$. In particular, the last estimate shows that $w\in BMO(B^{\varepsilon}_{\bar{r}},d_{\varepsilon},|\cdot|)$ as the ball $\widetilde{B}^{\varepsilon}_{\rho}(\hat{g})\subset B^{\varepsilon}_{\bar{r}}$ is an arbitrary ball in the metric space $(B^{\varepsilon}_{\bar{r}},d_{\varepsilon})$.
[…]
(4.22) $\left(\frac{1}{|B^{\varepsilon}_{\rho}|}\int_{B^{\varepsilon}_{\rho}}v^{\bar{\kappa}(p-1)}\right)^{\frac{1}{\bar{\kappa}}}\leq C\left(\frac{1}{|B^{\varepsilon}_{\rho}|}\int_{B^{\varepsilon}_{\rho}}v^{-\bar{\kappa}(p-1)}\right)^{-\frac{1}{\bar{\kappa}}}\,$,
where $C>0$ depends only on $Q$ and $\bar{\kappa}\in(0,\hat{k})\subset(0,1/Q)$ (recall that actually $\bar{r}=2$ here).
@@
In order to proceed in the proof we go back to (4.19) and use the Sobolev inequality (2.6). For $\hat{g}=0$ we take the cut-off function $\zeta$ as above with $\varepsilon/2\leq\bar{r}/4\leq\rho_{1}<\rho_{2}\leq\bar{r}$ and we observe that $B_{\rho_{1}}^{\varepsilon}\subset spt\,\zeta\subset B_{\rho_{2}}^{\varepsilon}$ with uniformly comparable volumes because of (2.4), moreover Lemma 2.8 yields $|B_{\rho_{1}}^{\varepsilon}|\leq C\rho_{2}^{Q}$ for a constant $C>0$ depending only on $Q$. Thus, from Jensen inequality together with (2.6) and (2.4) we infer for any $\kappa\in\mathbb{R}\setminus\{0,1\}$ and $\varepsilon/2\leq\bar{r}/4\leq\rho_{1}<\rho_{2}\leq\bar{r}$ that
(4.23) $\left(\frac{1}{|B^{\varepsilon}_{\rho_{1}}|}\int_{B^{\varepsilon}_{\rho_{1}}}\left(v^{(p-1)}\right)^{\frac{\kappa Q}{Q-1}}\right)^{\frac{Q-1}{|\kappa|Q}}$ $\leq\left(\frac{1}{|B^{\varepsilon}_{\rho_{1}}|}\int_{B^{\varepsilon}_{\rho_{1}}}\left(v^{\kappa(p-1)}\right)^{\frac{Q}{Q-p}}\right)^{\frac{Q-p}{|\kappa|Q}}$
$\leq\left(Cp^{*}\rho_{2}\right)^{p/|\kappa|}\left[\frac{1}{|\kappa-1|}+1\right]^{p/|\kappa|}\left(\frac{1}{|B^{\varepsilon}_{\rho_{2}}|}\int_{B^{\varepsilon}_{\rho_{2}}}\left(v^{(p-1)}\right)^{\kappa}|\nabla_{\varepsilon}\zeta|_{\varepsilon}^{p}\right)^{1/|\kappa|}$
$\leq\left(\frac{Cp^{*}\rho_{2}}{\rho_{2}-\rho_{1}}\right)^{Q/|\kappa|}\left[\frac{1}{|\kappa-1|}+1\right]^{Q/|\kappa|}\left(\frac{1}{|B^{\varepsilon}_{\rho_{2}}|}\int_{B^{\varepsilon}_{\rho_{2}}}\left(v^{(p-1)}\right)^{\kappa}\right)^{1/|\kappa|},$
where the constant $C>1$ depends only on $Q$.
Next we perform Moser iteration with exponents $\{\kappa_{i}\}_{i\geq 0}\subset(0,1)\cup(1,\infty)$. For integer $j\geq 0$ we set $\kappa_{j}^{*}:=\left(\frac{Q}{Q-1}\right)^{-(2j+1)/2}\to 0$ as $j\to\infty$ and we chose $\bar{j}$ so large that by construction $\bar{\kappa}:=\kappa^{*}_{\bar{j}}<\hat{\kappa}$. Next we define a strictly increasing sequence $\kappa_{i}:=\left(\frac{Q}{Q-1}\right)^{i}\bar{\kappa}$, hence $\bar{\kappa}=\kappa_{0}<\kappa_{\bar{j}}=\sqrt{\frac{Q-1}{Q}}<1<\sqrt{\frac{Q}{Q-1}}=\kappa_{\bar{j}+1}$ and by monotonicity and trivial manipulations we get $|\kappa_{i}-1|\geq\frac{1}{2Q}$ for any $i\geq 0$.
@@
Thus, for fixed $\varepsilon\leq r=\bar{r}/2$ we repeatedly use (4.23) with the choice $(\rho_{1},\rho_{2},\kappa)=(\tilde{\rho}_{i+1},\tilde{\rho}_{i},\kappa_{i})$, where $\tilde{\rho}_{i}=\bar{r}(\frac{1}{2}+2^{-i-1})$ for $i\geq 0$, so that in view of the lower bound $|\kappa_{i}-1|\geq\frac{1}{2Q}$ inequality(4.23) rewrites for all $i\geq 0$ as
(4.24) $\left(\frac{1}{|B^{\varepsilon}_{\tilde{\rho}_{i+1}}|}\int_{B^{\varepsilon}_{\tilde{\rho}_{i+1}}}\left(v^{p-1}\right)^{\kappa_{i+1}}\right)^{\frac{1}{|\kappa_{i+1}|}}\leq(Cp^{*})^{\frac{Q(i+2)}{|\kappa_{i}|}}\left(\frac{1}{|B^{\varepsilon}_{\tilde{\rho}_{i}}|}\int_{B_{\tilde{\rho}_{i}}^{\varepsilon}}\left(v^{p-1}\right)^{\kappa_{i}}\right)^{\frac{1}{|\kappa_{i}|}}\,\,$,
for a constant $C>1$ depending only on $Q$.
@@
(4.25) $\left(\underset{B^{\varepsilon}_{r}}{\rm{ess}\sup}\,\,v\right)^{p-1}=\|v\|^{p-1}_{L^{\infty}(B^{\varepsilon}_{r})}=\left\|v^{p-1}\right\|_{L^{\infty}(B^{\varepsilon}_{r})}\leq(Cp^{*})^{C^{\prime}}\left(\frac{1}{|B_{\bar{r}}^{\varepsilon}|}\int_{B_{\bar{r}}^{\varepsilon}}v^{\bar{\kappa}(p-1)}\right)^{\frac{1}{\bar{\kappa}}}\,$,
where $C>1$ depends on $Q$ and $C^{\prime}=\sum_{i=0}^{\infty}Q\frac{i+2}{|\kappa_{i}|}<\infty$ depends only on $Q$ (as it is the case for $\bar{k}$).
On the other hand, for fixed $\varepsilon\leq r=\bar{r}/2$ relying again on (4.23) we perform another Moser iteration to show that
(4.26) $\left(\underset{B^{\varepsilon}_{r}}{\rm{ess}\inf}\,\,v\right)^{-(p-1)}=\left\|v^{-(p-1)}\right\|_{L^{\infty}(B^{\varepsilon}_{r})}\leq(Cp^{*})^{C^{\prime}}\left(\frac{1}{|B^{\varepsilon}_{\bar{r}}|}\int_{B^{\varepsilon}_{\bar{r}}}v^{-\bar{\kappa}(p-1)}\right)^{1/\bar{\kappa}}\,$,
with $C$ and $C^{\prime}$ exactly as above, i.e., depending only on $Q$. Indeed, if for $i\geq 0$ we set $\tilde{\kappa}_{i}:=-\kappa_{i}\searrow$-$\infty$ as $i\to\infty$ then $\tilde{\kappa}_{0}=-\bar{\kappa}$ and $\{\tilde{\kappa}_{i}\}\subset(-\infty,0)$, hence we can apply (4.23) with the choice $(\rho_{1},\rho_{2},\kappa)=(\tilde{\rho}_{i+1},\tilde{\rho}_{i},\tilde{\kappa}_{i})$ and where still $\tilde{\rho}_{i}=\bar{r}(\frac{1}{2}+2^{-i-1})$ for $i\geq 0$. Thus, we get once more (4.24) for all $i\geq 0$ for $\tilde{\kappa}_{i}$ instead of $\kappa_{i}$. Moreover, as in the proof of (4.25), iterating (4.24) for these new set of exponents inequality (4.26) follows easily as $i\to\infty$ .
Finally, combining inequalities (4.25) and (4.26) with (4.22) for $\rho=\bar{r}$, the desired inequality (4.16) follows with $C>1$ and $C^{\prime}>0$ with the desired dependence on $Q$ only.